OpenStudy (anonymous):

Will someone please help on these three questions? For number 1 and 2 they have the same question but different problems. Choose an equation for the line that is through the given point for questions 1 and 2. 1)y=5x+10;(2,14) a)y=1/5x+4 b)y=-1/5x+4 c)y=5x-68 d)y=5x+4 2)y=1/5x-10; (15,-16) a)y=-5x+19 b)y=1/5x+91/5 c)y=1/5x-19 d)y=5x-19 5.choose the equation of a line that perpendicular to the given line and that passes through the given point. Ax-12y=2; (10,-1) a)y=3x+29 b)y=-1/3x+29 c)y=-3x+29 d)y=-1/3x+7

OpenStudy (anonymous):

Well the first one is 5x + 4 I just graphed it to get that and the second is -3x + 29 and I again graphed it :)

OpenStudy (anonymous):

sorry the third one is -3x + 29 I didn't really see the second one

OpenStudy (texaschic101):

1. y = 5x + 10 -- In y = mx + b form, the slope is in the m position, therefore, the slope of this line is 5. now use y = mx + b slope(m) = 5 (2,14)..x = 2 and y = 14 now we sub, and since we already know the slope(m), we need to find the y intercept(b) 14 = 5(2) + b 14 = 10 + b 14 - 10 = b 4 = b so your equation is : y = 5x + 4 ==================== y = 1/5x - 10 y = mx + b slope(m) = 1/5 (15,-16)..x = 15 and y = -16 time to sub and find the y intercept(b) -16 = 1/5(15) + b -16 = 3 + b -16 - 3 = b -19 = b so your equation is : y = 1/5x - 19 ======================= not sure about the 3rd one...sorry

OpenStudy (anonymous):

well there ya go I answered the third one he gave you the first and second ;)

OpenStudy (anonymous):

thanks y'all I really appreciate all y'alls help :)